A soda company want to stimulate sales in this economic climate by giving customers a chance to win a small prize for ever bottle of soda they buy. There is a 30% chance that a customer will find a winning icon at the bottom of the cap upon opening up a bottle of soda. The customer can then redeem that bottle cap for a small prize. Now, if yours truly buys a 6-pack of soda
My question below
what is the probability that I will win something, i.e., at least winning a single small prize?

Answer by stanbon(75887)

A soda company want to stimulate sales in this economic climate by giving customers a chance to win a small prize for ever bottle of soda they buy. There is a 30% chance that a customer will find a winning icon at the bottom of the cap upon opening up a bottle of soda. The customer can then redeem that bottle cap for a small prize. Now, if yours truly buys a 6-pack of soda
My question below
what is the probability that I will win something, i.e., at least winning a single small prize?
----
Probability a single bottle will NOT win a prize:: 0.7
----
P(none of 6 will win a prize) = 0.7^6 = 0.12
-------
P(at least one of the 6 is a winner) = 1-0.12 = 0.88
------------
Cheers,
Stan H.
-----------